Two practically elastic spheres collide. If their velocities before impact have same magnitude and opposite directions and one has a mass 3 times the other, what are their velocities after impact?
1
Expert's answer
2021-10-10T18:11:21-0400

Let the velocity of the spheres before collision u

And the mass of them is m and 3m

Let there speed after collision be "v_{1}" and "v_{2}"

Hence,

Velocity of approach = velocity of separation

"u_{1}-u_{2}=v_{2}-v_{1} \\

u-(-u)=v_{2}-v_{1} \\

v_{2}-v_{1}=2u \\

v_{2}=2u+v_{1}...(i)"

Now from the conservation of momentum

"m_{1}u_{1}+m_{2}u_{2}=m_{1}v_{1}+m_{2}v_{2} \\

mu+3m(-u)=mv_{1}+3mv_{2} \\

mu-3mu=mv_{1}+3mv_{2} \\

-2mu=mv_{1}+3mv_{2} \\

v_{1}+3v_{2}=-2u... (ii)"

Solve (i) and (ii)

"v_{1}+3(2u+v_{1})=-2u \\

v_{1}+6u+3v_{1}=-2u \\

v_{1}+3v_{1}=-2u-6u \\

4v_{1}=-8u \\

v_{1}=-2u"

Putting this in (i)

"v_{2}=2u+(-2u) \\

v_{2}=0"

hence the 1st sphere will come back with a speed twice of its initial and the 2nd sphere will attains at rest.